# MoE-FFD: Mixture of Experts for Generalized and Parameter-Efficient Face Forgery Detection

### v2

@@ 4.4 Ablation Experiments
4.4.4 MoE v.s. Multi-Experts.
While the proposed MoE learning scheme exhibits superior generalizability compared to using a single expert, it is still ambiguous whether the performance boosts stem from the designed MoE or the joint usage of experts. To further demonstrate the dynamic selection of optimal LoRA and Adapter experts by MoE, we compare MoE with Multi-Experts (Multi-E). Multi-E aggregates the features of all designed experts. The efficiency and the face forgery detection performance of MoE and Multi-E are reported in Table VIII. Thanks to the used gating mechanism that selectively activates the sparse experts, our MoE achieves a 1.28$\times$ speedup in training and 1.23$\times$ speedup in inference. Despite these efficiency gains, we interestingly find that MoE consistently outperforms Multi-E in terms of the cross-dataset detection performance. This suggests that naively aggregating multiple experts may not be the optimal strategy for face forgery detection. One potential explanation for this phenomenon is that using all feature extractor experts could suppress the most informative features while introducing noisy ones. The results in Table VIII further underscores the effectiveness of the proposed method in selecting the optimal expert for face forgery detection.
4.4.5 Effectiveness of MoE Loss.
The gating network often exhibits a tendency to consistently assign large weights to only a few experts [48], resulting in overfitting problems. To address this issue, we introduce an MoE loss component aimed at encouraging equal importance among all experts. This regularization also prevents the model from getting trapped in local optima. In this subsection, we examine the impacts of the proposed $L_{moe}$. Table IX presents the cross-dataset evaluation performance with different values of the $L_{moe}$ loss weight $\lambda$ in Eq. (5), where $\lambda=0$ represents no MoE loss applied. We observe that the use of $L_{moe}$ effectively mitigates the gate overfitting problem and consistently boosts model generalizability. Furthermore, the model achieves the best generalizability with $\lambda=1$.
4.5 Visualization Results
TABLE IX: Effectiveness of loss components.
@@ 5 Conclusions and Future Works
In this paper, we introduced MoE-FFD, a generalized yet parameter-efficient method for detecting face forgeries. By incorporating external lightweight LoRA and Adapter layers with the frozen ViT backbone, our framework adaptly acquired forgery-specific knowledge with minimal activated parameters. This approach not only harnesses the expressiveness of transformers but also capitalizes on the local forgery priors with customized adapters, contributing to enhanced detection performance. Through dynamic expert selection within both LoRA and Adapter layers, our MoE design further enhances the model’s generalizability and robustness. Extensive experiments consistently demonstrated MoE-FFD’s superiority in face forgery detection across diverse datasets, manipulation types, and perturbation scenarios. Moreover, MoE-FFD serves as a parameter-efficient detector and can seamlessly adapt to various ViT backbones, facilitating its deployment and fine-tuning in real-world applications. Last but not least, comprehensive ablation studies demonstrated the effectiveness of our designed LoRA layers and Convpass Adapter layers, and the MoE learning scheme indeed helped the model to search the optimal forgery experts. We anticipate that MoE-FFD will inspire future advancements in face forgery detection, particularly in bolstering generalization and efficiency.
While our MoE-FFD framework achieves generalized and robust performance in image-level face forgery detection, adapting our network design and learning scheme to video-level Deepfake detection is worth considering. Moreover, detecting audio-visual content forgery is essential in real world applications, multimodal Deepfake detection opens an important research path forward. We envision that our proposed MoE structure and the PEFT strategy should be still effective in these tasks. This path can be explored in our future work.

### v3

@@ 4.5 Visualization Results
TABLE XII: Effectiveness of loss components.
@@ 5 Conclusions and Future Works
In this paper, we introduced MoE-FFD, a generalized yet parameter-efficient method for detecting face forgeries. By incorporating external lightweight LoRA and Adapter layers with the frozen ViT backbone, our framework adaptively acquired forgery-specific knowledge with minimal activated parameters. This approach not only harnesses the expressiveness of transformers but also capitalizes on the local forgery priors via customized adapters, thereby significantly enhancing the detection performance. Through dynamic expert selection within both LoRA and Adapter layers, our MoE design further boosts the model’s generalizability and robustness. Extensive experiments consistently demonstrated MoE-FFD’s superiority in face forgery detection across diverse datasets, manipulation types, and perturbation scenarios. Moreover, MoE-FFD serves as a parameter-efficient detector and can seamlessly adapt to various ViT backbones, facilitating its deployment and fine-tuning in real-world applications. Last but not least, comprehensive ablation studies demonstrated the effectiveness of our designed LoRA layers and Convpass Adapter layers, and the MoE learning scheme indeed helped the model to search the optimal forgery experts. We anticipate that MoE-FFD will inspire future advancements in face forgery detection, particularly in bolstering generalization and efficiency.
While our MoE-FFD framework achieves generalized and robust performance in image-level face forgery detection, extending our network design and learning scheme to video-level Deepfake detection represents an important future research direction. Moreover, detecting audio-visual content forgery is essential in real world applications, multimodal Deepfake detection opens an important research path forward. We envision that our proposed MoE structure and the PEFT strategy should be still effective in these tasks. This path can be explored in our future work.